OpenStudy (anonymous):

What is the equation of the line that passes through (1, 2) and is parallel to the line whose equation is 4x + y - 1 = 0? 4x - y - 6 = 0 4x + y - 6 = 0 4x + y + 6 = 0

OpenStudy (anonymous):

It must be a type of equation like 4x+y+C=0 Plug in (1,2) 4*1+2+C=0 C+6=0 C=-6 4x+y-6=0